OpenStudy (anonymous):

A circle is represented by the equation below: (x + 2)2 + (y − 4)2 = 225 Which statement is true?

OpenStudy (anonymous):

statements are?

OpenStudy (anonymous):

The circle is centered at (−2, 4) and has a radius of 15. The circle is centered at (2, −4) and has a diameter of 15. The circle is centered at (2, −4) and has a radius of 15. The circle is centered at (−2, 4) and has a diameter of 15.

OpenStudy (anonymous):

@study100 @nikato

OpenStudy (campbell_st):

the general form is \[(x - h)^2 + (y - k)^2 = r^2\] (h, k) is the centre... and r is the radius so match it to your equation... and the radius is \[\sqrt{225} = r\]

OpenStudy (anonymous):

i am confused how to olug in the numbers @campbell_st

OpenStudy (anonymous):

sorry I meant center****

OpenStudy (campbell_st):

ok... so in the bracket with x, what is the number... in your question...?

OpenStudy (anonymous):

center (h,k)

OpenStudy (campbell_st):

so you have \[(x + 2^2).. and (x - h)^2\] any thoughts on the value of h

OpenStudy (anonymous):

2? @campbell_st

OpenStudy (campbell_st):

well its 2 = -h..... so whats the value of h...?

OpenStudy (campbell_st):

same thing with k \[(y - 4)^2 ...and....((y - k)^2\] find the value of k

OpenStudy (campbell_st):

well if 2 = -h then h = -2 can you try finding the k value...?

OpenStudy (anonymous):

value of k is 4? @campbell_st

OpenStudy (campbell_st):

thats correct... so the centre is (h, k) so what do you think the centre is..?

OpenStudy (campbell_st):

so your centre is (-2, 4) and you need to find the radius r\[\sqrt{225} = r\] and that will be all you need

OpenStudy (anonymous):

I have no idea. @campbell_st

OpenStudy (campbell_st):

here is a scientific calculator that will help you find \[\sqrt{225}\] http://web2.0calc.com/

OpenStudy (anonymous):

15 @campbell_st

OpenStudy (campbell_st):

thats correct... now match your findings to the answers... centre (-2, 4) radius 15

OpenStudy (anonymous):

Choice A :) Thank u @campbell_st

OpenStudy (campbell_st):

great effort...